How would i simplify 4/7 - (3/2)?
Found 2 solutions by juliane1, penguin18:
Answer by juliane1(5)

2^2/7 - 3/2
2(2^20 + 7 (-3)/7*2
2*4 -7*3/7*2
8-21/7*2
-13/7*2
13/14

Answer by penguin18(7)

Use common denominators (least common multiple of both numbers, in this case 14.)
4/7*2/2=8/14
3/2*7/7=21/14
8/14-21/14=-13/14